What is the distance between y-intercepts of the graph of x+8=2(y+3)^2?
Answer by ewatrrr(24785)

x+8=2(y+3)^2 y-intercepts (x = 0)
4 = (y+3)^2
± 2 = y+3
-3 ±2 = y
y is -1, -5
distance between (0,-1) and (0,-5) = 4